Solve the folling 2 x 2 systems using substitutions
y= 1/2x + 10
y= 5x +1

Answer by checkley75(3666)

Y=5X+1
5X+1=1/2X+10
5X-1/2X=10-1
4.5X=9
X=9/4.5
X=2 ANSWER. NOW SUBSTITUTE 2 FOR X & SOLVE FOR Y
Y=5*2+1
Y=10+1
Y=11 ANSWER.
PROOF
11=1/2*2+10
11=1+10
11=11
------------------------------------------------
WE ARE HERE TO HELP SOLVE PROBLEMS NOT TO DO YOUR HOMEWORK.
SAVE US A LOT OF EXTRA WORK BY SUBMITTING ONLY A FEW (1 OR 2) OF EACH TYPE OF PROBLEM -- PLEASE. WE HAVE PLENTY OF PEOPLE LOOKING FOR A SINGLE PROBLEM SOLUTION BUT YOU ARE TAKING UP TOO MUCH SPACE SO WE CAN'T GET TO THE OTHERS WHO ALSO HAVE NEEDS.